OpenStudy (anonymous):

16) find an equation of the line having the given slope and containing the given points m= -6,(6,0) use format y=mx+b

OpenStudy (anonymous):

use the almighty point slope formula \[y-y_1=m(x-x_1)\]

OpenStudy (anonymous):

with \[m=-6, x_1=6,y_1=0\] to get \[y-0=-6(x-6)\] \[y=-6x+12\]